Louisiana State University Health Sciences Center Administration & Finance
 

Friday, July 20, 2007

NOTICE: On Saturday, July 21st from 10:00 AM to 7:00 PM there will be no condenser water or domestic water available at the Clinical Sciences Research Building .  Also, the steam to the CSRB will be shut down at the same time.  This means that the CSRB environmental rooms will lose temperature control and there will be no steam for the autoclaves in the CSRB.  Also, no toilets will be operable and no water will be available at the faucets.  Please refrain from using toilets and opening faucets during the outage.  The outage is needed to move piping so that equipment can be raised up to bring it above the flood level experienced in 2005.

Dental School : 

  • The installation of the ceiling tile on the 8th floor of the Clinic Building is nearly complete.
    • Landis continues to clean on the 2nd through the 7th floors of the Clinic Building.
    • Ceiling grid has been completely installed on the 2nd through the 4th floors of the Administration Building.
    • Ceiling tiles are being installed throughout the 2nd floor of the Administration Building.
    • The installation of the light fixtures in the Library is nearing completion.
  • The remediation contractor is prepping the railing and doors in Stairwell B in the Clinic building for painting.  
    • Zimmer is re-cleaning the coils on AHUs 7A and 7C in the Clinic Building. 
    • The units will be re-tested when the cleaning is complete.
  • The termination of the new network cables continued in the Administration Building.
    • The final steps in constructing of the new network closets are nearing completion.
    • Work includes door installation and the completion of the new walls on several floors in the Clinic building.
    • Network and phone drops have been installed in the temporary work areas on the 1st floor of the Clinic Building.
  • All of the new walls in the Central Sterilization unit have been framed.
    • The remaining drywall is waiting for the installation of the door frames.  installation of the door frames. 
    • The fire hose cabinet piping was removed to allow for installation of the new drywall. 
  • Carpet installation is nearly complete in the Clinic Building with the exception of a few rooms where the ceiling work is not yet finished. 
    • This will protect the new carpet from being damaged by any debris created by the tile installation.
  • Work on the Change Order to create temporary work spaces for Supply Chain Management, Facility Services and Central Supply on the 1st floor of the Clinic Building has begun.
    • Materials are being staged.
    • Electricians were on site and started work,
  • The installation of the new cooling tower work continues with a shutdown of the system this evening.
    • This will allow the contractor to pipe the new cooling tower into the condenser water system and complete all of the electrical and control work. 
  • The installation of the doors on the 1st floors required by the State Fire Marshall is nearly competed.
    • All of the door frames have been installed. 
    • The door hardware is expected next week and this will complete the project.

1542 Tulane Ave :

  • Ordes connected panel PL-4 and is terminating wires in sub-panels PLL-6 & PLL7.
    • In addition, Ordes installed a Current Transformer Meter (CT Meter) for panel PL-4.
  • Bell Roofing is on site setting up for the removal of the shed on the 9th floor roof. The actual removal is scheduled for Sunday.

Student Lot Repaving:

  • The contractor poured the new island in Lot #3 and backfilled the grade on the corner of Perdido & Bertrand Street.
    • The contractor patched the asphalt along the curb and the corner on Perdido Street.
  • The contractor backfilled the grade around the new sitting area in Lot #2 and patched the asphalt along the curb and the corner on Perdido Street .